# TacMan-Turbo: Proactive Tactile Control for Robust and Efficient Articulated Object Manipulation

### v2

@@ 4.4 Simulation Results
Both methods achieve perfect $100\%$ manipulation success rates across all $200$ test objects spanning $6$ categories, as shown in Table 1. This consistent performance demonstrates the fundamental effectiveness of tactile-informed approaches for articulated object manipulation. For comprehensive comparisons between tactile-informed approaches and other manipulation strategies, we refer readers to the detailed analysis in Zhao et al[73].
Figure 5: Time efficiency comparison (logarithmic scale, seconds). Evaluation of task completion times reveals TacMan-Turbo’s superior efficiency compared to Tac-Man across all tested articulation types. Our method consistently completes manipulation tasks in significantly less time, with performance advantages particularly pronounced for complex articulations. Statistical analysis confirms these improvements are highly significant (****$p<0.0001$) throughout all joint categories. Box plots display time distributions with medians (center lines), interquartile ranges (boxes), and distribution extent (whiskers at $1.5\timesIQR$). BC-n denotes Bézier curves of order n.
@@
Figure 6: Analysis of action efficiency across manipulation strategies. Comprehensive evaluation reveals TacMan-Turbo’s superior motion efficiency through three complementary measures. Temporal profiles highlight TacMan-Turbo’s continuous productive motion vs. Tac-Man’s alternating execution-recovery cycles. Mean efficiency measurements confirm TacMan-Turbo’s consistently higher percentage of productive motion across all object categories. Relative standard deviation analysis demonstrates TacMan-Turbo’s more stable and predictable performance. Statistical tests across 200 objects spanning six joint categories show highly significant improvements (**** $p<0.0001$) in both efficiency magnitude and consistency. BC-n represents Bézier curves of order n; box plots follow conventions established in Fig. 5.
@@
Figure 7: Analysis of motion smoothness through jerk characteristics. Motion smoothness evaluation reveals TacMan-Turbo maintains consistently lower jerk magnitudes by eliminating the disruptive transition phases inherent to Tac-Man’s approach. The temporal comparison shows TacMan-Turbo’s continuous motion profile contrasted with Tac-Man’s oscillatory pattern. Statistical analysis across 200 test objects confirms TacMan-Turbo achieves dramatic smoothness gains with standard joint types (prismatic and revolute), while improvements become more moderate with complex Bézier trajectories where challenging pose predictions occasionally necessitate corrective motions. Nevertheless, significant overall improvement (**** $p<0.0001$) in motion quality is observed across all categories. Note: BC-n denotes Bézier curves of order n; box plots follow conventions established in Fig. 5.
@@
While maintaining identical success rates, TacMan-Turbo delivers dramatic improvements in execution speed, as detailed in Fig. 5. Average task completion times for TacMan-Turbo remain consistently around $10~\mathrm{s}$ across all six object categories, compared to Tac-Man’s typical requirement of over $100~\mathrm{s}$. Statistical analysis across all $200$ test objects confirms this substantial difference between TacMan-Turbo ($M=7.6~\mathrm{s},SD=2.3~\mathrm{s}$) and Tac-Man ($M=250.7~\mathrm{s},SD=194.4~\mathrm{s}$), $t(199)=17.7,p=2.1\times 10^{-52}$.
@@
The architectural differences between the two methods are most evident in their action efficiency profiles. As illustrated in Fig. 6(a), Tac-Man’s reactive approach requires dedicated recovery periods that, while necessary for correcting contact deviations, do not directly advance the manipulation task. In contrast, TacMan-Turbo’s proactive approach ensures that virtually all actions contribute directly to task progression, eliminating these non-productive recovery phases.
Our quantitative analysis examines both mean efficiency and the Relative Standard Deviation (RSD)—the ratio of standard deviation to mean—across all trials. Results presented in Figs. 6(b) and 6(c) demonstrate TacMan-Turbo’s consistently higher average efficiency with minimal variation, even when handling complex Bézier curve trajectories. Statistical analysis across all $200$ test objects confirms significant improvements in average action efficiency between TacMan-Turbo ($M=0.239~\%,SD=0.063~\%$) and Tac-Man ($M=0.012~\%,SD=0.011~\%$), $t(199)=52.63,p=8.9\times 10^{-133}$. Additionally, efficiency stability shows marked improvement, with TacMan-Turbo ($M=1.0,SD=1.0$) demonstrating significantly lower fluctuation than Tac-Man ($M=8.8,SD=6.9$), $t(199)=15.9,p=7.0\times 10^{-38}$.
@@
Trajectory smoothness analysis reveals fundamental behavioral differences between the two approaches. As shown in Fig. 7(a), TacMan-Turbo demonstrates more consistent and significantly lower jerk variations compared to Tac-Man. This improvement stems directly from TacMan-Turbo’s proactive strategy, which eliminates the frequent transitions between execution and recovery stages characteristic of Tac-Man. In the reactive approach, compensatory actions to address contact deviations typically require abrupt accelerations and decelerations, resulting in higher jerk measurements.
Comprehensive testing across all $200$ objects, summarized in Fig. 7(b), reveals varying degrees of improvement across object categories. TacMan-Turbo shows the most dramatic smoothness gains with standard prismatic and revolute-joint objects, with more moderate improvements for Bézier-curve trajectories. This performance variation correlates with motion predictability—standard joint types follow consistent patterns that facilitate accurate prediction, while Bézier curves present more variable motion patterns that occasionally lead to mismatches between predicted and actual handle poses. Nevertheless, overall smoothness analysis confirms statistically significant improvement between TacMan-Turbo ($M=109.82~\mathrm{s}^{-3},SD=147.71~\mathrm{s}^{-3}$) and Tac-Man ($M=4921.44~\mathrm{s}^{-3},SD=3722.49~\mathrm{s}^{-3}$), $t(199)=18.27,p=9.11\times 10^{-45}$.
5 Real-World Experiments
While simulation studies establish TacMan-Turbo’s theoretical advantages, real-world deployment introduces critical challenges absent in simulation: sensor noise, execution uncertainty, and physical constraints. These factors can significantly impact control performance, necessitating comprehensive physical validation. This section presents our real-world experimental evaluation in four complementary parts: experimental setup (Section 5.1), testing protocols (Section 5.2), performance analysis (Section 5.3), and validation on everyday objects (Section 5.4)—including categories not represented in standard articulated object datasets.
@@ 5 Real-World Experiments
5.1 Real-World Experimental Setup
@@
Manipulating articulated objects with varied handle paths requires full 6-DoF movement capability in Cartesian space. Our platform employs a 7-DoF Kinova Gen3 robotic arm paired with a Robotiq 2F-85 gripper, providing the necessary workspace coverage and manipulation flexibility. To enable precise contact sensing—crucial for our tactile-based approach—we modified the gripper by replacing its standard pads with two GelSight-type tactile sensors. This integrated system, illustrated in Fig. 8, combines motion versatility with tactile feedback. Overall, the tactile sensor updates contact information at $120\text{\,}\mathrm{Hz}$, while the offset velocity is computed at $50\text{\,}\mathrm{Hz}$ and transmitted to Kinova’s low-level controller, which operates at $1000\text{\,}\mathrm{Hz}$.
@@
Figure 10: Tactile-informed pose estimation of in-hand objects. Precise tracking of object movements during manipulation is demonstrated through tactile feedback across three different objects subjected to five distinct manipulations: initial state, pressing down, dragging upward, and two types of twisting motions. Contact points (marked with purple dots) are tracked to compute relative pose transformations using Eq. 9. The colored grids display marker displacement patterns while grayscale images reveal normal force distribution. Quantitative relative translations ($m$) and rotations (axis-angle representation) below each image pair confirm accurate capture of subtle pose changes. Visualization arrows (scaled 7× for clarity) illustrate movement direction and magnitude, showing consistent correspondence between applied manipulations and detected transformations across varied object geometries.
@@
Figure 11: Real-world experiment results. Comparative analysis across three articulated objects (drawer, microwave, vise) demonstrates successful manipulation by both methods with clear performance differences. Each row presents a complete manipulation sequence showing initial stable contact (left), manipulation stages (middle), and corresponding 3D trajectories (right). Tactile feedback visualizations display contact force distributions through marker displacement patterns (colored grids) and normal force intensity (grayscale images). While both approaches successfully complete all tasks, TacMan-Turbo consistently delivers more direct trajectories with fewer corrective movements, resulting in reduced completion time, higher action efficiency, and smoother motion profiles. The trajectory plots particularly highlight TacMan-Turbo’s more continuous paths (red) vs. Tac-Man’s characteristic stop-and-go patterns (blue)
@@
Figure 12: Tests on various common household objects. Our TacMan-Turbo approach successfully generalizes to diverse everyday objects without requiring additional tuning. The system effectively manipulates: (a) a battery charging dock requiring precise device extraction, (b) a power switch with discrete on/off state transitions, (c) a coffee maker with rotational lid mechanism, and (d) a sewing machine featuring complex mechanical linkages. Before (top) and after (bottom) images demonstrate successful manipulation for each object. These tests validate TacMan-Turbo’s ability to adapt to varied kinematic constraints and interaction requirements found in common household items, confirming its practical applicability beyond the primary test objects.
@@
Accurate contact information between the object and robot is essential for TacMan-Turbo’s operation. Our implementation utilizes two GelSight-type tactile sensors [67], each featuring an $8\times 8$ grid of black markers (Fig. 8). These markers enable precise computation of the relative transformation matrix between gripper and handle through point correspondence matching, as formalized in Eq. 9.
To validate the sensing system’s reliability, we conducted controlled perturbation tests with in-hand objects (Fig. 10). The computed transformations consistently aligned with applied perturbation directions, confirming the system’s ability to accurately estimate relative pose changes—a capability fundamental to the prediction framework described in Section 3.3.
The identification of contact points requires careful parameter calibration to ensure accurate selection within the contact region (Eq. 2). In our implementation, we dynamically adjust the contact threshold $\epsilon$ in Eq. 2 during each trial to the smallest value that yields exactly 9 contact points in the initial contact set $\mathcal{C}_{i}$. This adjustment satisfies the mathematical constraints specified in Eqs. 11 and 12 given a maximum of 8 markers along any single line in our setting, ensuring robust contact deviation computation.
@@
Following our simulation methodology, we evaluate TacMan-Turbo across three kinematically distinct categories: a drawer with prismatic joints (Fig. 9(a)), a microwave oven with revolute joints (Fig. 9(b)), and a vise featuring a helical handle path (Fig. 9(c)). This selection enables systematic assessment across fundamental joint types while incorporating real-world mechanical complexities such as friction variations and mechanical play.
5.2 Real-World Experiment Protocol
@@
Each experimental trial began with identical initialization conditions: manually established stable grasps on object handles (illustrated in the left column of Fig. 11), ensuring consistent starting contact states. We maintained controlled experimental parameters by testing both methods on the same physical objects with identical initial interaction directions, eliminating potential confounding variables. For the grasping force, since a larger force ensures the non-slipping assumption, we used the lesser of two values: the gripper’s maximum force or the tactile sensor’s force tolerance. In the case of the selected Robotiq 2F-85 gripper, its maximum force is significantly lower than the tactile sensor’s capacity. Therefore, we consistently applied the gripper’s maximum force.
@@
5.3 Real-World Experiment Results
Our experimental results, presented in Fig. 11, provide comprehensive evidence of both methods’ manipulation capabilities while revealing significant performance distinctions between their approaches.
@@ 5.3 Real-World Experiment Results
Both TacMan-Turbo and Tac-Man successfully manipulated all test objects shown in Fig. 9, confirming the fundamental effectiveness of tactile-based strategies for articulated object manipulation. Each method maintained stable contact throughout the manipulation process and completed all assigned tasks. The resulting motion trajectories (middle-right column, Fig. 11) closely followed expected kinematic patterns: linear motion for the prismatic-joint drawer, circular arcs for the revolute-joint microwave door, and extended helical paths for the vise handle. Minor deviations primarily stemmed from manufacturing tolerances inherent in real-world objects.
@@
Despite similar success rates, time efficiency analysis revealed dramatic performance differences between the two approaches. As captured in the middle-left column of Fig. 11, TacMan-Turbo completed manipulation tasks substantially faster, while Tac-Man had barely initiated object movement at the same point in time. This efficiency gap stemmed from fundamental differences in operational strategies evident in detailed trajectory analysis (right column, Fig. 11).
@@
5.4 Additional Tests on Everyday Objects
@@
A battery charging dock (Fig. 12(a)) exemplifies objects with linear motion paths, where precise trajectory following is essential for secure device removal and to prevent potential damage. A power switch (Fig. 12(b)) represents safety-critical mechanisms with discrete state transitions, requiring reliable manipulation despite high resistance forces. A coffee maker (Fig. 12(c)) involves rotational motion, demanding adaptive handling to ensure smooth operation. Lastly, a sewing machine (Fig. 12(d)) introduces intricate motion patterns with varying resistance caused by its mechanical linkages and operational states.
@@
6 Discussions
6.1 Comparison with Tac-Man
Both TacMan-Turbo and Tac-Man leverage robot-object contact as the state representation. This approach ensures robustness without requiring pre-defined object kinematics for articulated object manipulation. The effectiveness of this strategy is comprehensively demonstrated through simulation results (Table 1) and validated in real-world experiments (Fig. 11).

### v3

@@ IV-D Simulation Results
As shown in Tab.˜I, the pre-planned method achieves considerable success rates across all tests with known kinematics. However, its performance degrades substantially under parameter estimation uncertainty, highlighting its sensitivity to modeling errors. Impedance control demonstrates effectiveness in handling moderate directional deviations without requiring known kinematics or tactile feedback. Nevertheless, its adaptive capacity becomes limited when large discrepancies exist between the planned and actual manipulation directions. The learning-based method performs well on objects with visual appearances similar to those encountered during training. However, its success rate drops to nearly zero when confronted with unfamiliar objects, revealing the generalization limitations inherent to current learning-based approaches.
[…]
Figure 5: Time efficiency comparison (logarithmic scale, seconds). Evaluation of task completion times reveals TacMan-Turbo’s superior efficiency compared to Tac-Man across all tested articulation types. Our method consistently completes manipulation tasks in significantly less time, with performance advantages particularly pronounced for complex articulations. Statistical analysis confirms these improvements are highly significant (****$p<0.0001$) throughout all joint categories. BC-n denotes Bézier curves of order n; box plots follow conventions established in Fig.˜4.
@@
Figure 6: Analysis of action efficiency across manipulation strategies. The results reveal TacMan-Turbo’s superior motion efficiency through three complementary measures. Temporal profiles highlight TacMan-Turbo’s continuous productive motion vs. Tac-Man’s alternating execution-recovery cycles. Mean efficiency measurements confirm TacMan-Turbo’s consistently higher percentage of productive motion across all object categories. Relative standard deviation analysis demonstrates TacMan-Turbo’s more stable and predictable performance. Statistical tests across 200 objects spanning six joint categories show highly significant improvements (**** $p<0.0001$) in both efficiency magnitude and consistency. BC-n represents Bézier curves of order n; box plots follow conventions established in Fig.˜4.
@@
Figure 7: Analysis of motion smoothness through jerk characteristics. Motion smoothness evaluation reveals TacMan-Turbo maintains consistently lower jerk magnitudes by eliminating the disruptive transition phases inherent to Tac-Man’s approach. The temporal comparison shows TacMan-Turbo’s continuous motion profile contrasted with Tac-Man’s oscillatory pattern. Statistical analysis across 200 test objects confirms TacMan-Turbo achieves dramatic smoothness gains with standard joint types (prismatic and revolute), while improvements become more moderate with complex Bézier trajectories where challenging pose predictions occasionally necessitate corrective motions. Nevertheless, significant overall improvement (**** $p<0.0001$) in motion quality is observed across all categories. Note: BC-n denotes Bézier curves of order n; box plots follow conventions established in Fig.˜4.
@@
[…]
@@
The architectural differences between the two methods are most evident in their action efficiency profiles. As illustrated in Fig.˜6(a), Tac-Man’s reactive approach requires dedicated recovery periods that, while necessary for correcting contact deviations, do not directly advance the manipulation task. In contrast, TacMan-Turbo’s proactive approach ensures that virtually all actions contribute directly to task progression, eliminating these non-productive recovery phases.
[…]
@@
Trajectory smoothness analysis reveals fundamental behavioral differences between the two approaches. As shown in Fig.˜7(a), TacMan-Turbo demonstrates more consistent and significantly lower jerk variations compared to Tac-Man. This improvement stems directly from TacMan-Turbo’s proactive strategy, which eliminates the frequent transitions between execution and recovery stages characteristic of Tac-Man. In the reactive approach, compensatory actions to address contact deviations typically require abrupt accelerations and decelerations, resulting in higher jerk measurements.
[…]
V Real-World Experiments
While simulation studies establish TacMan-Turbo’s theoretical advantages, real-world deployment introduces critical challenges absent in simulation: sensor noise, execution uncertainty, and physical constraints. These factors can significantly impact control performance, necessitating comprehensive physical validation. This section presents our real-world experimental evaluation in four complementary parts: experimental setup (Sec.˜V-A), testing protocols (Sec.˜V-B), performance analysis (Sec.˜V-C), and validation on everyday objects (Sec.˜V-D)—including categories not represented in standard articulated object datasets.
@@ V Real-World Experiments
V-A Real-World Experimental Setup
@@
Manipulating articulated objects with varied handle paths requires full 6-DoF movement capability in Cartesian space. Our platform employs a 7-DoF Kinova Gen3 robotic arm paired with a Robotiq 2F-85 gripper, providing the necessary workspace coverage and manipulation flexibility. To enable precise contact sensing—crucial for our tactile-based approach—we modified the gripper by replacing its standard pads with two GelSight-type tactile sensors. This integrated system, illustrated in Fig.˜8, combines motion versatility with tactile feedback. The tactile sensor’s pose relative to the robot wrist is determined from the robot’s mechanical model. The internal camera of the GelSight-type sensor is then calibrated using OpenCV’s built-in routines: we estimate the camera intrinsics and lens distortion coefficients by imaging laser-cut black markers that serve as calibration targets.
[…]
@@
Figure 10: Tactile-informed pose estimation of in-hand objects. Precise tracking of object movements during manipulation is demonstrated through tactile feedback across three different objects subjected to five distinct manipulations: initial state, pressing down, dragging upward, and two types of twisting motions. Contact points (marked with purple dots) are tracked to compute relative pose transformations using Eq.˜8. The colored grids display marker displacement patterns while the grayscale images reveal normal force distribution. Quantitative relative translations ($m$) and rotations (axis-angle representation) below each image pair confirm accurate capture of subtle pose changes. Visualization arrows (scaled 7× for clarity) illustrate movement direction and magnitude, showing consistent correspondence between applied manipulations and detected transformations across varied object geometries.
@@
Figure 11: Real-world experiment results. Comparative analysis across three articulated objects (drawer, microwave, vise) demonstrates successful manipulation by both methods with clear performance differences. Each row presents a complete manipulation sequence showing initial stable contact (left), manipulation stages (middle), and corresponding 3D trajectories (right). Tactile feedback visualizations display contact force distributions through marker displacement patterns (colored grids) and normal force intensity (grayscale images). While both approaches successfully complete all tasks, TacMan-Turbo consistently delivers more direct trajectories, resulting in reduced completion time and smoother motion profiles. Detailed comparisons are available in the Supplementary Video S3.
@@
Figure 12: Tests on various common household objects. Our TacMan-Turbo approach successfully generalizes to diverse everyday objects without requiring additional tuning. The system effectively manipulates: (a) a battery charging dock requiring precise device extraction, (b) a power switch with discrete on/off state transitions, (c) a coffee maker with a rotational lid mechanism, and (d) a sewing machine featuring complex mechanical linkages. Before (top) and after (bottom) images demonstrate successful manipulation for each object. These tests validate TacMan-Turbo’s ability to adapt to varied kinematic constraints and interaction requirements found in common household items, confirming its practical applicability beyond the primary test objects. Detailed manipulation processes are available in the Supplementary Video S4.
@@
Accurate contact information between the object and robot is essential for TacMan-Turbo’s operation. Our implementation utilizes two GelSight-type tactile sensors [64], each featuring an $8\times 8$ grid of black markers (Fig.˜8). These markers enable precise computation of the relative transformation matrix between gripper and handle through point correspondence matching, as formalized in Eq.˜8.
To validate the sensing system’s reliability, we conducted controlled perturbation tests with in-hand objects (Fig.˜10). The computed transformations consistently aligned with applied perturbation directions, confirming the system’s ability to accurately estimate relative pose changes—a capability fundamental to the prediction framework described in Sec.˜III-C.
Selecting contact points requires careful parameter tuning to ensure accurate identification within the contact region; see Eq.˜2. As noted by Zhao et al. [1], the transition zone between contact and non-contact regions can exhibit substantial normal deformation due to the soft sensor’s compliance, yet it does not reliably reflect changes in handle position. This makes a fixed threshold $\epsilon$ ineffective.
To mitigate this, we adapt $\epsilon$ in Eq.˜2 during each trial, choosing the smallest value that yields exactly nine contact points in the initial contact set $\mathcal{C}_{i}$. Given that our setup places at most eight markers on any single line, this choice satisfies the constraints in Eqs.˜10 and 11 by preventing collinearity, thereby enabling robust computation of contact deviations.
@@
Following our simulation methodology, we evaluate TacMan-Turbo across three kinematically distinct categories: a drawer with prismatic joints (Fig.˜9(a)), a microwave oven with revolute joints (Fig.˜9(b)), and a vise featuring a helical handle path (Fig.˜9(c)). This selection enables systematic assessment across fundamental joint types while incorporating real-world mechanical complexities such as friction variations and mechanical play.
V-B Real-World Experiment Protocol
@@
Each experimental trial began with identical initialization conditions: manually established stable grasps on object handles (illustrated in the left column of Fig.˜11), ensuring consistent starting contact states. We maintained controlled experimental parameters by testing both methods on the same physical objects with identical initial interaction directions, eliminating potential confounding variables. For the grasping force, since a larger force ensures the non-slipping assumption, we used the lesser of two values: the gripper’s maximum force or the tactile sensor’s force tolerance. In the case of the selected Robotiq 2F-85 gripper, its maximum force is significantly lower than the tactile sensor’s capacity. Therefore, we consistently applied the gripper’s maximum force.
@@
V-C Real-World Experiment Results
Our experimental results, presented in Fig.˜11, provide evidence of both methods’ manipulation capabilities while revealing significant performance distinctions between their approaches.
@@ V-C Real-World Experiment Results
Both TacMan-Turbo and Tac-Man successfully manipulated all test objects shown in Fig.˜9, confirming the fundamental effectiveness of tactile-based strategies for articulated object manipulation. Each method maintained stable contact throughout the manipulation process and completed all assigned tasks. The resulting motion trajectories (middle-right column, Fig.˜11) closely followed expected kinematic patterns: linear motion for the prismatic-joint drawer, circular arcs for the revolute-joint microwave door, and extended helical paths for the vise handle. Minor deviations primarily stemmed from manufacturing tolerances inherent in real-world objects.
@@
Despite similar success rates, time efficiency analysis revealed dramatic performance differences between the two approaches. As captured in the middle-left column of Fig.˜11, TacMan-Turbo completed manipulation tasks substantially faster, while Tac-Man had barely initiated object movement at the same point in time. This efficiency gap stemmed from fundamental differences in operational strategies evident in detailed trajectory analysis (right column, Fig.˜11).
@@
V-D Additional Tests on Everyday Objects
@@
A battery charging dock (Fig.˜12(a)) exemplifies objects with linear motion paths, where precise trajectory following is essential for secure device removal and to prevent potential damage. A power switch (Fig.˜12(b)) represents safety-critical mechanisms with discrete state transitions, requiring reliable manipulation despite high resistance forces. A coffee maker (Fig.˜12(c)) involves rotational motion, demanding adaptive handling to ensure smooth operation. Lastly, a sewing machine (Fig.˜12(d)) introduces intricate motion patterns with varying resistance caused by its mechanical linkages and operational states.
@@
VI Discussions
VI-A Comparison with Tac-Man
Both TacMan-Turbo and Tac-Man leverage robot-object contact as the state representation. This approach ensures robustness without requiring pre-defined object kinematics for articulated object manipulation. The effectiveness of this strategy is comprehensively demonstrated through simulation results (Tab.˜I) and validated in real-world experiments (Fig.˜11).